# Robustness-aware Automatic Prompt Optimization

### v2

@@ E.1 Number of iterations
In our adversarial training process, we set the number of iterations to five, as we observed that the prompts optimized by the LLM achieved optimal performance at the fifth iteration. To illustrate this, we used the text summarization task, a computationally intensive text generation task, as an example. We tested the performance of prompts generated across six iterations (from 1 to 6), and the results are presented in the  Figure 15.
Figure 15: The relationship between the number of iterations and the performance of the prompt generated by BATprompt.
@@
For each iteration, we evaluate the prompts across different perturbation types and calculate their average performance. The results show a consistent improvement in prompt performance during the initial rounds, with the optimal performance observed around the fourth or fifth iteration. This trend suggests that the LLM effectively refines the prompts through iterative gradient-guided optimization and semantic space exploration, progressively approaching the optimal solution. However, beyond the fifth iteration, the performance of the prompts havea slightly declines. In order to speed up the inference experiment and reduce the cost, we choose the number of iterations to be 5
E.2 Performance on the original dataset
In this section, we evaluate the performance metrics of prompts generated by BATprompt when applied to tasks on unperturbed datasets. This assessment demonstrates that our method is not only robust to perturbed data but also has greate performance on unperturbed data. The results for the language understanding tasks are illustrated in the  Figure 4. P1 and P2 represent different prompts generated by BATprompt for two types of perturbations. As shown in the  Figure 4, BATprompt achieves the best performance on several datasets (e.g., SST-5 and TREC). For other tasks, even when it does not outperform all methods, its performance is comparable to the best results (e.g., SST-2 and CR).
Similarly, the results for text generation tasks are presented in  Figure 16, with the left figure illustrating the text summarization task and the right figure depicting the text simplification task. The experimental findings indicate that BATprompt also achieves the best performance on text generation tasks. Notably, in the text summarization task, its metrics are significantly higher than those of the second-best method. In conclusion, the prompts generated by BATprompt not only maintain robustness on perturbed datasets but also deliver strong performance on original, unperturbed datasets.
@@ E.2 Performance on the original dataset
Figure 16: The performance of the prompts generated by BATprompt on undisturbed datasets of text simplification task and text summarization task.
Appendix F Future Work
@@
Exceptions Explore: During our testing of the effects of different types of perturbations on task pairs in  section 3, some intriguing observations emerged. While most perturbations negatively impacted task performance, certain perturbations surprisingly enhanced task execution. As shown in the Table 16, on the SST-5, AG’s News, and Asset datasets, some perturbations appeared to unlock the latent potential of the large model, enabling it to perform better on the tasks.
@@
We hypothesize that applying these perturbations to the input introduces diversity, encouraging the LLM to engage its reasoning abilities rather than relying on specific representations. Additionally, such perturbations may influence the model’s attention distribution, helping to resolve ambiguities in the input and ultimately enhancing task performance. This presents a highly valuable avenue for exploration. We could investigate a fixed perturbation strategy that consistently enhances the performance of LLMs when processing text inputs.
@@
Appendix G Optimal Prompts
We publish the prompts that are optimal on different tasks after BATprompt generation and the prompts for Manual Instruction and Natural Instruction as baseline. Including language understanding( Table 17 and  Table 18), text summarization( Table 19), and text simplification( Table 20).
@@ Appendix G Optimal Prompts
BATprompt(P1) $\cdot$ Decide if the text expresses a ’negative’ or ’positive’ sentiment without taking into account any extra details, even if there are mistakes in spelling and typos compared to the original text. $\cdot$ For this assignment, you will receive sentences that have been altered from movie reviews. Your job is to determine whether the sentiment of the sentence is positive or negative and classify it accordingly as either "positive" or "negative.
BATprompt(P2) $\cdot$ Conduct sentiment analysis by categorizing the sentiment of a sentence as ńegativeór ṕositive.́ Output only the sentiment label with no other information. $\cdot$ For this assignment, you will be provided with sentences taken from movie reviews. Your goal is to determine whether each sentence conveys a positive or negative sentiment by classifying it as either "positive" or "negative." The reviews may discuss specific characters, actions, critiques, relationships between characters, performances, or negative aspects of the works.
@@
BATprompt(P1) $\cdot$ Analyze the sentiment of the modified text as either ’negative’ or ’positive’, disregarding any intentional spelling and grammar mistakes, and provide only the corresponding label as the result. $\cdot$ For this assignment, you will receive original movie reviews. Your goal is to determine if a sentence has a positive sentiment by classifying it as "positive," or if it has a negative sentiment by classifying it as "negative.
BATprompt(P2) $\cdot$ Complete a Sentiment Classification task by analyzing a modified text with enhanced details and extra information. Provide a sentiment label of either ńegativeór ṕositive,́ and only submit the label. $\cdot$ For this assignment, you will receive modified sentences from movie reviews. Your goal is to categorize a sentence as "positive" if it expresses positive sentiment or as "negative" if it expresses negative sentiment, regardless of any changes made to the text such as including negative adjectives or conveying different meanings using synonyms.
@@
BATprompt(P1) $\cdot$ Analyze the sentiment of the revised text given, ignoring any mistakes in spelling, typos, or edits made. Categorize the sentiment as either ’positive’ or ’negative’ without any additional details. $\cdot$ In this assignment, you will be given altered versions of movie reviews. Your goal is to classify each sentence as either positive for examining deeper themes and emotions, or negative for focusing solely on surface-level aspects of the films.
BATprompt(P2) $\cdot$ Complete the Sentiment Classification task by assigning a sentiment label of either ’negative’ or ’positive’ to the provided shortened sentence and return the label only. $\cdot$ For this task, you will be given distorted sentences from movie reviews. Your objective is to identify whether a sentence conveys a negative sentiment and label it as ńegative,́ or if it conveys a positive sentiment and label it as ṕositive.́
@@
BATprompt(P1) $\cdot$ Perform a Sentiment Classification task by assigning a sentiment label from [’terrible’,’bad’, ’okay’, ’good’, ’great’] to the modified sentence. Only include the label, no extra information needed. $\cdot$ During this activity, you will be presented with sentences from movie reviews that have been modified to provide more general and subjective opinions on the movies. Your task remains the same: classify each review into one of the categories - terrible, bad, okay, good, or great - based on its sentiment and tone.
BATprompt(P2) $\cdot$ Perform Sentiment Classification. Given the altered text, assign a sentiment label from [’terrible’, ’bad’, ’okay’, ’good’, ’great’]. Return the label only. $\cdot$ In this task, you are given sentences from movie reviews. Based on the given review, classify it to one of the five classes: (1) great, (2) good, (3) okay, (4) bad, and (5) terrible.
Table 17: Specific prompt of Manual Instruction(baseline), Natural Instruction, BATprompt in P1 and BATprompt in P2 in language understanding task.
Dataset Method Content
@@
BATprompt(P1) $\cdot$ Complete a News Classification assignment. Select a category from [Ẃorld,́ Śports,́ B́usiness,́ T́ech]́ for a news item that may have different spellings and minor grammar mistakes compared to the original text. Only provide the label without extra information. $\cdot$ For this assignment, you will be provided with a news article. Your goal is to categorize the article into one of four topics: "World," "Sports," "Business," or "Tech" based on its main focus. If the article discusses trends in the U.S. stock market or gaming news, select either the "Business" or "Tech" category depending on the emphasis. If the article includes nonsensical or random words that do not make sense, choose the "Other" category.
BATprompt(P2) $\cdot$ Classify news by assigning a label from the options [’World’, ’Sports’, ’Business’, ’Tech’] to each news item. Only provide the assigned label, without any extra text. $\cdot$ For this task, you will receive a news article and your objective is to classify it under one of the four categories: "World", "Sports", "Business", "Tech" based on the articleś main subject. If you are uncertain about the category, please select the most appropriate option. The Revised text includes extra details on different subjects and individuals like Tommy Tuberville and Major League Soccer, which are not referenced in the Original text. Additionally, the Revised text offers more precise information and examples, such as the mishandling of the conferenceś top position and the difficulties faced by Major League Soccer.
@@
BATprompt(P1) $\cdot$ Determine the appropriate category for the text by providing only the label without any extra details. Choose from Description, Entity, Expression, Human, Location, or Number. $\cdot$ Identify the appropriate category for the given text by selecting from "Description", "Entity", "Expression", "Human", "Location", and "Number" depending on the context.
BATprompt(P2) $\cdot$ Complete a Question Classification activity where you are provided with a shortened version of the original question without key words or phrases, and categorize it into one of the following labels: Description, Entity, Expression, Human, Location, or Number. Provide only the assigned label as the output. $\cdot$ Decide on the correct category for the text provided. Select from "Description", "Entity", "Expression", "Human", "Location", and "Number".
Table 18: Specific prompt of Manual Instruction(baseline), Natural Instruction, BATprompt in P1 and BATprompt in P2 in language understanding task.
@@
BATprompt(P1) Please give a concise overview of the main idea communicated in the text.
BATprompt(P2) Identify the main idea or central theme of the text.
Table 19: Manual Instructions as the baseline and instructions with best performance generated by BATprompt (either P1 or P2) on Xsum.
@@
BATprompt(P1) Rephrase the text using easier words without including any additional details.
BATprompt(P2) Rephrase the text using more straightforward language and clearer wording to enhance understanding.
Table 20: Manual Instructions as the baseline and instructions with best performance generated by BATprompt (either P1 or P2) on Asset.

### v3

@@ D.1 Number of iterations
In our optimization process, we set the number of iterations to five, as we observed that the prompts optimized by the LLMs achieved optimal performance at the fifth iteration. To illustrate this, we used the text summarization task, a computationally intensive text generation task, as an example. We tested the performance of prompts generated across six iterations (from 1 to 6), and the results are presented in the Experiments.
@@
For each iteration, we evaluate the prompts across different perturbation types and calculate their average performance. The results show a consistent improvement in prompt performance during the initial rounds, with the optimal performance observed around the fourth or fifth iteration. This trend suggests that the LLMs effectively refines the prompts through iterative gradient-guided optimization and semantic space exploration, progressively approaching the optimal solution. However, beyond the fifth iteration, the performance of the prompts have a slightly declines. In order to speed up the inference experiment and reduce the cost, we choose the number of iterations to be 5
D.2 Performance on the original dataset
@@
Figure 12: The performance of the prompts generated by PGO on undisturbed datasets of classification task.
In this section, we evaluate the performance metrics of prompts generated by PGO when applied to tasks on unperturbed datasets. This assessment demonstrates that our method is not only robust to perturbed data but also has greate performance on unperturbed data. The results for the classification tasks are illustrated in the Figure. 12. P1 and P2 represent different prompts generated by PGO for two types of perturbations. As shown in the Figure. 12, PGO achieves the best performance on several datasets (e.g., SST-5 and TREC). For other tasks, even when it does not outperform all methods, its performance is comparable to the best results (e.g., SST-2 and CR).
Similarly, the results for text generation tasks are presented in Figure. 13, with the left figure illustrating the text summarization task and the right figure depicting the text simplification task. The experimental findings indicate that PGO also achieves the best performance on text generation tasks. Notably, in the text summarization task, its metrics are significantly higher than those of the second-best method. In conclusion, the prompts generated by PGO not only maintain robustness on perturbed datasets but also deliver strong performance on original, unperturbed datasets.
Figure 13: The performance of the prompts generated by PGO on undisturbed datasets of text simplification task and text summarization task.
Appendix E Future Work
@@
Exceptions Explore: During our testing of the effects of different types of perturbations on task pairs, some intriguing observations emerged. While most perturbations negatively impacted task performance, certain perturbations surprisingly enhanced task execution. As shown in the Table. 23, on the SST-5, AG’s News, and Asset datasets, some perturbations appeared to unlock the latent potential of the large model, enabling it to perform better on the tasks.
@@
We hypothesize that applying these perturbations to the input introduces diversity, encouraging the LLMs to engage its reasoning abilities rather than relying on specific representations. Additionally, such perturbations may influence the model’s attention distribution, helping to resolve ambiguities in the input and ultimately enhancing task performance. This presents a highly valuable avenue for exploration. We could investigate a fixed perturbation strategy that consistently enhances the performance of LLMs when processing text inputs.
@@
Appendix F Optimal Prompts
Iteration Prompts: We optimized our prompt over five rounds, and the table below presents the final results for each iteration. We use the summarization task as examples. Since our approach aims to make LLMs better suited for the task while exploring the perturbed semantic space during optimization, the resulting prompts differ significantly from the original version. The results in the Table. 28.
Prompts for contrasting methods: We publish the prompts that are optimal on different tasks after PGO generation and the prompts for Manual Instruction and Natural Instruction as baseline. Including classification(Table. 24 and Table. 25), text summarization(Table. 26), and text simplification(Table. 27). From classification prompts, we can observe that for DA prompt, we observe that, due to the absence of explicit guidance, its prompt words and MI structure remain largely unchanged. This indicates that it fails to fully explore the semantic space to generate optimal prompts. For Evoprompt, it lacks targeted strategies to address perturbations. As a result, it can’t help LLMs resist perturbations well. Our prompts not only explore a broader semantic space but also incorporate specific terms to mitigate the effects of perturbation. From language generation prompts, we can see that for Evoprompt contains excessive redundant information, such as "so readers can comprehend the important concepts and essential information.". This may affect the LLMs understanding. for simple iterative prompt, It suffers from the same problem as DA’s prompts, which exhibits high similarity with MI, suggesting that the LLMs fails to explore a broader semantic space for generating more effective prompts. Our prompts actively navigate a richer semantic space, leading to more diverse and informative outputs.
@@ Appendix F Optimal Prompts
Evoprompt Examine the movie reviews and classify them as either positive or negative.
Data-Augmentation Please assign a sentiment label (’negative’ or ’positive’) to the given sentence for the Sentiment Classification task. Submit only the label without any extra information. This task is expected to have better performance.
PGO(P1) $\cdot$ Decide if the text expresses a ’negative’ or ’positive’ sentiment without taking into account any extra details, even if there are mistakes in spelling and typos compared to the original text. $\cdot$ For this assignment, you will receive sentences that have been altered from movie reviews. Your job is to determine whether the sentiment of the sentence is positive or negative and classify it accordingly as either "positive" or "negative.
PGO(P2) $\cdot$ Conduct sentiment analysis by categorizing the sentiment of a sentence as ńegativeór ṕositive.́ Output only the sentiment label with no other information. $\cdot$ For this assignment, you will be provided with sentences taken from movie reviews. Your goal is to determine whether each sentence conveys a positive or negative sentiment by classifying it as either "positive" or "negative." The reviews may discuss specific characters, actions, critiques, relationships between characters, performances, or negative aspects of the works.
@@
Evoprompt Analyze customer reviews and categorize each sentence as either ’positive’ or ’negative’.
Data-Augmentation Please complete the Sentiment Classification task by providing a sentiment label (either ’negative’ or ’positive’) for the given sentence. Only return the assigned label without any additional text. This task should have improved performance.
PGO(P1) $\cdot$ Analyze the sentiment of the modified text as either ’negative’ or ’positive’, disregarding any intentional spelling and grammar mistakes, and provide only the corresponding label as the result. $\cdot$ For this assignment, you will receive original movie reviews. Your goal is to determine if a sentence has a positive sentiment by classifying it as "positive," or if it has a negative sentiment by classifying it as "negative.
PGO(P2) $\cdot$ Complete a Sentiment Classification task by analyzing a modified text with enhanced details and extra information. Provide a sentiment label of either ńegativeór ṕositive,́ and only submit the label. $\cdot$ For this assignment, you will receive modified sentences from movie reviews. Your goal is to categorize a sentence as "positive" if it expresses positive sentiment or as "negative" if it expresses negative sentiment, regardless of any changes made to the text such as including negative adjectives or conveying different meanings using synonyms.
@@
Evoprompt Identify if a movie review is positive or negative by accurately categorizing each input-output pair into either ’positive’ or ’negative’.
Data-Augmentation Please provide a sentiment label (’negative’ or ’positive’) for the given sentence in the Sentiment Classification task. Submit only the label with no additional details. The task is anticipated to show improved performance.
PGO(P1) $\cdot$ Analyze the sentiment of the revised text given, ignoring any mistakes in spelling, typos, or edits made. Categorize the sentiment as either ’positive’ or ’negative’ without any additional details. $\cdot$ In this assignment, you will be given altered versions of movie reviews. Your goal is to classify each sentence as either positive for examining deeper themes and emotions, or negative for focusing solely on surface-level aspects of the films.
PGO(P2) $\cdot$ Complete the Sentiment Classification task by assigning a sentiment label of either ’negative’ or ’positive’ to the provided shortened sentence and return the label only. $\cdot$ For this task, you will be given distorted sentences from movie reviews. Your objective is to identify whether a sentence conveys a negative sentiment and label it as ńegative,́ or if it conveys a positive sentiment and label it as ṕositive.́
Table 24: Specific prompt of Manual Instruction(baseline), Natural Instruction, PGO in P1 and PGO in P2 in classification task.
Dataset Method Content
@@
Evoprompt Have your friend evaluate the movie they had just seen and provide a summary opinion (e.g. terrible, bad, okay, good, or great) to determine the sentiment of the movie review.
Data-Augmentation Perform the sentiment classification task by assigning a sentiment label of either ’terrible’, ’bad’, ’okay’, ’good’, or ’great’ to the given sentence. Return only the label without any additional text.
PGO(P1) $\cdot$ Perform a Sentiment Classification task by assigning a sentiment label from [’terrible’,’bad’, ’okay’, ’good’, ’great’] to the modified sentence. Only include the label, no extra information needed. $\cdot$ During this activity, you will be presented with sentences from movie reviews that have been modified to provide more general and subjective opinions on the movies. Your task remains the same: classify each review into one of the categories - terrible, bad, okay, good, or great - based on its sentiment and tone.
PGO(P2) $\cdot$ Perform Sentiment Classification. Given the altered text, assign a sentiment label from [’terrible’, ’bad’, ’okay’, ’good’, ’great’]. Return the label only. $\cdot$ In this task, you are given sentences from movie reviews. Based on the given review, classify it to one of the five classes: (1) great, (2) good, (3) okay, (4) bad, and (5) terrible.
@@
Evoprompt Please classify news articles into categories such as World, Sports, Business, or Tech. Output only the label without any additional text.
Data-Augmentation Please complete the Sentiment Classification task by providing a sentiment label (either ’negative’ or ’positive’) for the given sentence. Only return the assigned label without any additional text. This task should have improved performance.
PGO(P1) $\cdot$ Complete a News Classification assignment. Select a category from [Ẃorld,́ Śports,́ B́usiness,́ T́ech]́ for a news item that may have different spellings and minor grammar mistakes compared to the original text. Only provide the label without extra information. $\cdot$ For this assignment, you will be provided with a news article. Your goal is to categorize the article into one of four topics: "World," "Sports," "Business," or "Tech" based on its main focus. If the article discusses trends in the U.S. stock market or gaming news, select either the "Business" or "Tech" category depending on the emphasis. If the article includes nonsensical or random words that do not make sense, choose the "Other" category.
PGO(P2) $\cdot$ Classify news by assigning a label from the options [’World’, ’Sports’, ’Business’, ’Tech’] to each news item. Only provide the assigned label, without any extra text. $\cdot$ For this task, you will receive a news article and your objective is to classify it under one of the four categories: "World", "Sports", "Business", "Tech" based on the articleś main subject. If you are uncertain about the category, please select the most appropriate option. The Revised text includes extra details on different subjects and individuals like Tommy Tuberville and Major League Soccer, which are not referenced in the Original text. Additionally, the Revised text offers more precise information and examples, such as the mishandling of the conferenceś top position and the difficulties faced by Major League Soccer.
@@
Evoprompt Recognize the inputs (explanations, entities, or humans) and provide the suitable outputs (numbers, descriptions, or entities) to answer the questions in a way that is understandable for non-native English speakers.
Data-Augmentation Please label the questions using one of the options: ’Description’, ’Entity’, ’Expression’, ’Human’, ’Location’, ’Number’. Only provide the label, no extra text needed.
PGO(P1) $\cdot$ Determine the appropriate category for the text by providing only the label without any extra details. Choose from Description, Entity, Expression, Human, Location, or Number. $\cdot$ Identify the appropriate category for the given text by selecting from "Description", "Entity", "Expression", "Human", "Location", and "Number" depending on the context.
PGO(P2) $\cdot$ Complete a Question Classification activity where you are provided with a shortened version of the original question without key words or phrases, and categorize it into one of the following labels: Description, Entity, Expression, Human, Location, or Number. Provide only the assigned label as the output. $\cdot$ Decide on the correct category for the text provided. Select from "Description", "Entity", "Expression", "Human", "Location", and "Number".
Table 25: Specific prompt of Manual Instruction(baseline), Natural Instruction, PGO in P1 and PGO in P2 in classification task.
@@
Evoprompt Reduce the core by reading or listening carefully to identify the main ideas and key points, so readers can comprehend the important concepts and essential information.
Data-Augmentation How can you summarize that with better performance?
Simple-Iteration How can the performance of the summary task be enhanced?
PGO(P1) Identify the main idea or central theme of the text.
PGO(P2) Determine the primary focus or central message of the text.
Table 26: Manual Instructions as the baseline and instructions with best performance generated by PGO (either P1 or P2) on Xsum.
@@
Evoprompt Rewrite the given sentence to make it more accessible and understandable for both native and non-native English speakers.
Data Augmentation Rewrite the text for improved performance.
Simple-Iteration Rewrite the text to improve its overall effectiveness.
PGO(P1) Rephrase the text using easier words without including any additional details.
PGO(P2) Rephrase the text using more straightforward language and clearer wording to enhance understanding.
Table 27: Manual Instructions as the baseline and instructions with best performance generated by PGO (either P1 or P2) on Asset.
Method Content
I1 Can you provide a brief overview of the key points discussed in the text?
I2 Summarize the main ideas presented in the text in a clear and succinct manner.
I3 Condense the key points and important facts from the text into a concise summary that captures the main ideas effectively
I4 Create a brief and compelling summary that encapsulates the main ideas and significance of the text
I5 Determine the primary focus or central message of the text.
Table 28: Prompts in different Iterations in PGO.